OpenStudy (anonymous):

convert x^2 +y^2= a^2 to a polar equation

OpenStudy (anonymous):

\(x^2+y^2=a^2\) is the equation of a circle with radius \(a\). When converting to polar, you have \(x^2+y^2=r^2\), so you get \[r^2=a^2\] Since \(r>0\), when you take the square root, you disregard the negative root, leaving you with: \[r=a\]

OpenStudy (anonymous):

but to make it a polar equation don't you have to get rid of the a?

OpenStudy (anonymous):

No, \(a\) is just a constant.

OpenStudy (anonymous):

okay, thanks!!

OpenStudy (anonymous):

how would you do r=theta?

OpenStudy (anonymous):

\(r=\sqrt{x^2+y^2}\) and \(\tan\theta=\dfrac{y}{x}\).

OpenStudy (anonymous):

make that a rectangular equation

OpenStudy (anonymous):

so would it be \[x^{2}+y ^{2}=\frac{ y }{ x }\]

OpenStudy (anonymous):

Close, you have \(\tan\theta=\dfrac{y}{x}\), so \(\theta=\arctan\dfrac{y}{x}\).

OpenStudy (anonymous):

the first equation is only valid if it's r^2

OpenStudy (anonymous):

\[r^2=x^2+y^2~~\Rightarrow~~r=\sqrt{x^2+y^2}\] disregarding the negative root, since \(r\) is necessarily non-negative.

OpenStudy (anonymous):

oh!!! thanks!

OpenStudy (anonymous):

you're welcome